2.   The joint density function for random variables X, Y, and Z is

f (x,y,z) = {Cxyz, if 0 ≤ x ≤ 2, 0 ≤ y ≤ 2, 0 ≤ z ≤ 2

0, otherwise


a)   Find the value of C.

b) Find P(X ≤ 1,Y ≤ 1, Z ≤ 1).


1
Expert's answer
2022-01-24T17:02:07-0500

Here, "f(x,y,z)" is a probability density function.

So,

a) "\int_0^2\int_0^2\int_0^2 Cxyzdzdydx=1"

"\implies" "\int_0^2\int_0^2\left\{\frac{Cxyz²}{2}\right\}_0^2dydx=1"


"\implies" "\int_0^2\int_0^22Cxydydx=1"


"\implies" "\int_0^2\left\{Cxy²\right\}_0^2dx=1"


"\implies" "\int_0^24Cxdx=1"


"\implies" "\left\{\frac{4Cx²}{2}\right\}_0^2=1"


"\implies" "\left\{2Cx²\right\}_0^2=1"


"\implies" "2C(2)\u00b2-2C(0)\u00b2=1"


"\implies" "8C=1"


"\implies" "C=\frac{1}{8}"


b) "P(X\u22641,Y\u22641,Z\u22641)"

"=\int_0^1\int_0^1\int_0^1\frac{xyz}{8}dzdydx"


"=\int_0^1\int_0^1\left\{\frac{xyz²}{16}\right\}_0^1dydx"


"=\int_0^1\int_0^1\frac{xy}{16}dydx"


"=\int_0^1\left\{\frac{xy²}{32}\right\}_0^1dx"


"=\int_0^1\frac{x}{32}dx"


"=\left\{\frac{x²}{64}\right\}_0^1=\frac{1}{64}"